A thermally insulated closed copper vessel contains water at 15°C, when the vessel is shaken vigorously for 15 minutes temperature rises to 17°C. The mass of the vessel is 100g and that of water is 200g. Neglect any thermal expansion. How much heat is transferred to the liquid vessel system? How much work has been done on this system? How much is the increase in the internal energy of the system? The specific heat of water is 4200 J/kg-K and for copper is 420 J/kg-K  

Grade:upto college level

1 Answers

Nirmal Singh.
10 years ago
T1 = 15 degree C , T2 = 17 degree C, m1 (Vessel) = 100g = 100x100 -2 Kg, m2 = 200g = 200x100 -2 Kg, , Specific heat of copper = s1 = 420 j/kg-K, specific heat of water s2 =4200 j/kg-K
(a) The heat transferred to the liquid vessel system is 0. The internal heat is shared in between the
vessel and water.
(b) Work done on the system = Heat produced
dw = m1 s1 (T2-T1) + m2 s2(T2 - T1)
dw = 100 × 10–3 × 420 × 2 + 200 × 10–3 × 4200 × 2 = 84 + 84 × 20 = 84 × 21 = 1764 J.
(c) dQ = 0, dU = – dw = 1764. [since dw = –ve work done on the system]
Thanks & Regards,
Nirmal Singh
Askiitians faculty